OpenStudy (anonymous):

Help please..."A laptop advertisement states that internal parts operate at 55° Celsius plus or minus 15°. Write an absolute value equation to represent the situation. Then solve the equation and write it as a compound inequality. State the meaning of the solution."

OpenStudy (anonymous):

so \(55\) is in the middle, and you can be \(15\) degrees higher, or \(15\) degrees lower what numbers are those?

OpenStudy (anonymous):

Um..I'm not sure what you mean by what numbers are those..

OpenStudy (anonymous):

what is 15 degrees more than 55?

OpenStudy (anonymous):

70??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is 15 degrees less than 55 degrees?

OpenStudy (anonymous):

40 ; so \[40<55<70?\]

OpenStudy (anonymous):

close

OpenStudy (anonymous):

it is always the case that \(40<55<70\) what you need is a variable to represent the permissible temperature your teacher probably uses \(x\) so if you put \(x\) as the possible temps, then the biggest \(x\) can be is 70 and the smallest is 40 replace the 55 in your above inequality by \(x\) and let me see what you get

OpenStudy (anonymous):

\[40<x <70\] so that's the inequality part?

OpenStudy (anonymous):

yes, or you might want to write \[40\leq x\leq 70\] so be more precise

OpenStudy (anonymous):

Got it :) How would I write an absolute value equation with this??

OpenStudy (anonymous):

you need a variable for that one as well again we use \(x\) as the possible temp and translate directly in to math the distance between \(x\) and \(55\) is written as an absolute value by \(|x-55|\)

OpenStudy (anonymous):

you want that distance to be less than or equal to 15 so write that let me know what you get

OpenStudy (anonymous):

Ohh ok so, \[|x-55|\le15\]

OpenStudy (anonymous):

you found the sound

OpenStudy (anonymous):

Yay! So I would solve this inequality right?

OpenStudy (anonymous):

guess what you get if you solve it?

OpenStudy (anonymous):

A medal?

OpenStudy (anonymous):

lol i already gave you a medal no you get the answer to the second question, which you answered first

OpenStudy (anonymous):

Right right I knew that lol... Ohh so I basically answered this backwards..

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\[|x-55|\le 15\iff -15\leq x-55\leq 15\iff 40\leq x\leq 70\]

OpenStudy (anonymous):

just two different ways of saying the same thing

OpenStudy (anonymous):

Thank you a ton for your help! I was going crazy trying to solve this problem.

OpenStudy (anonymous):

yw